### v1

@@ 1 Introduction
In practice, however, this approach is very unstable in general: if the window $g$ is well-localized in either time or frequency, which is the case for meaningful windows, this will be reflected in its ambiguity function $\mathcal{A}g$. More precisely, if, for example, $g$ decays fast in time, then $\mathcal{A}g(x,y)$ will also decay rapidly in the $x-$direction. The same is true for fast decay of $g$ in frequency, yielding a corresponding decay of $\mathcal{A}g$ in $y$.
Note that for many applications, it is desirable to work with a window that is well localized in time and frequency because, in this case, the STFT yields a representation of the signal that is better localized in time and frequency, i.e., it can better describe which frequencies occurred (roughly) at what time (with the limit being the uncertainty principle). Hence, even if $\mathcal{A}g$ is nonzero everywhere on $\mathbb{R}^{2}$, its magnitude typically decays very fast, making recovery of $\mathcal{A}f$ via Equation (3) highly unstable. Take for example the Gauss window $\varphi$. Its ambiguity function is a two-dimensional Gauss function, i.e.,
@@
The goal of this paper is to propose algorithms that can employ the convenient and explicit inversion formula (3) in a way that is numerically stable. Existing reconstruction algorithms are much more involved: the most prominent algorithms employed for a wide range of phase retrieval problems are iterative projection algorithms, such as the Gerchberg–Saxton algorithm [15] and Fienup’s algorithm [14], Averaged Alternating Reflections [7] and its variants [11, 13, 21, 20, 10]. A more recent popular iterative scheme is Wirtinger Flow [8]. There also exist methods based on convex relaxation, such as PhaseLift [9], which is based on the idea of lifting the problem to higher dimensions and then recovering a low-rank solution of a linear problem instead of solving the non-linear phase retrieval problem. In practice, however, these algorithms have not been adopted as the iterative methods mentioned before are computationally much more efficient.
Another method, specific to STFT phase retrieval, is phase gradient heap integration (PGHI), which recovers the phase of $V_{g}f$ point by point along a grid on the time-frequency plane [22]. By far, we have not described the entire literature on phase retrieval algorithms here. For a thorough description, we refer to [12].
In this paper, we propose a different approach to STFT phase retrieval. The idea is to use the simple and explicit inversion formula (3) but with measurements with respect to several windows instead of one. This way, we can give theoretical guarantees for stability (cf. Section 3).
It is important to emphasize that not only does the specific reconstruction formula (3) exhibit poor stability properties when used with a single window function, but Gabor phase retrieval itself (from a single window function) is inherently unstable [4]. The stability properties can be linked to the connectedness of the Gabor transform magnitude as a function in the time-frequency plane [18]. Stability can be restored when global recovery of the phase is relaxed and one only aims at recovering the phase on each connected component individually [3]. In this work, we intend to take a different route: instead of drawing measurements from one single window $g,$ we take the STFT magnitudes from several different windows. This way, we overcome the trade-off between stability and disconnectedness of the measurements. We aim at stably recovering the signal globally, instead of on each connected component individually (the latter being inherent to the single window setup). The central question we address in this work is the following:
@@
The rest of the paper is organized to address this key question. In Section 2, we propose two multi-window algorithms for STFT phase retrieval — one utilizing fractionally Fourier transformed dilated Gaussian windows, and the other employing Hermite functions of varying degrees. Section 3 establishes mathematical results on the improved inverse problem stability provided by these two multi-window approaches. In Section 4, we present several numerical examples, and conclusions are drawn in Section 5.
@@ 2 Multi-window approaches for STFT phase retrieval
In particular, we propose two architectures for $\left\{g_{j}\right\}_{j=1}^{N}$. The first is based on fractional Fourier transforms of dilated Gaussians, while the second is based on Hermite functions of different order. These choices are based on properties of the ambiguity functions of these windows. Before we describe the two measurement systems, we first record some preliminary facts that will be employed.
@@ 2.1 Dilated Gauss windows under the fractional Fourier transform
for some fixed dilation factor $a>1$ and some discretization parameter $N$ such that $\{\alpha_{j}\}_{j=1}^{N}$ is an equidistant discretization of the interval $[0,2\pi]$. In view of Equation (5), this choice of windows results in ambiguity functions $\mathcal{A}\phi_{j}$ that have an elliptical exponential decay stretched by a factor $a$ along the axis $(\cos\alpha_{j},\sin\alpha_{j}).$ The superposition of these “essential supports” is shaped in the form of a daffodil and approaches a disc shape by taking $N$ larger, which is also illustrated in Figure 1.
@@
Figure 1: The absolute value of the sum of multiple fractional Fourier transformed (dilated) Gauss windows plotted in the log scale. (a) $\mathcal{A}\varphi^{1}$, where $\varphi^{1}$ is the standard Gauss; (b) $\mathcal{A}\varphi^{2}$ for the dilated Gauss with $a=2$; (c) $\sum_{j=1}^{N}\mathcal{A}(\mathcal{F}_{\tilde{\alpha}_{j}}\varphi^{2})$ with $N=8$; (d) $\mathcal{A}\varphi^{10}$ for the dilated Gauss with $a=10$; (e) $\sum_{j=1}^{N}\mathcal{A}(\mathcal{F}_{{\alpha}_{j}}\varphi^{10})$ with $N=4$; (f) $\sum_{j=1}^{N}\mathcal{A}(\mathcal{F}_{{\alpha}_{j}}\varphi^{10})$ with $N=40$.
@@
The daffodil-shaped geometry shown in Figure 1e is a very simple construction and the reconstruction can be further simplified as follows. A simple calculation yields that the ambiguity functions of $\phi_{j}$ are real-valued and positive. Here, we use the notation $R_{\alpha_{j}}(x,y)=(x(\alpha_{j}),y(\alpha_{j}))$:
@@
through pointwise division in (6). We propose Algorithm 1 for the multi-window inversion using fractionally Fourier transformed Gaussian windows at various angles.
Algorithm 1 A Multiscale STFT Phase Retrieval Algorithm Based on the Fractionally Fourier Transformed Gaussian Windows.
@@
A drawback of this setup is that many windows are required to achieve a good support coverage of the time-frequency plane by their ambiguity functions, such as the disc shown in Figure 1f. Next, we propose a more technical construction, but with the advantage of requiring less data in the sense that fewer windows are necessary for a good support coverage.
@@ 2.2 Hermite windows
For this second architecture, the set of windows will consist of Hermite (also called Hermite–Gaussian) functions of increasing order. Their ambiguity functions have rotational symmetry with a number of circular root sets related to the order of the Hermite function; see Figures 2a, 2b, and 2c. The idea of our approach is to recover $\mathcal{A}f$ as if one was “peeling an onion”: for each Hermite function, we define an annulus on which $\mathcal{A}f$ can be retrieved stably. The union of these annuli results in a large disc in the time-frequency plane. In what follows, we will make this procedure more precise. For this, we start with the definition of the Hermite function $h_{n}$ of order $n$ [1]:
@@
The set of Laguerre functions $\{\mathfrak{L}_{n}(z)\}_{n\in\mathbb{N}}$ forms an orthogonal basis of $L^{2}(\mathbb{R})$ and
@@
The $n-$th Laguerre polynomial, and hence also the $n-$th Laguerre function, has exactly $n$ (strictly positive, pairwise different) roots. By Equation (7), $\mathcal{A}h_{n}$ is rotation-invariant. Hence, the root set of $\mathcal{A}h_{n}$ consists of $n$ disjoint circles, centered at the origin $(0,0)$ and thus has Lebesgue measure zero.
When solving the STFT phase retrieval problem with the help of (2), the stability of this recovery will be dictated by the absolute value of the ambiguity function: to recover the ambiguity function of the unknown signal via Equation (2) requires the division by the ambiguity function of the window. Therefore, this process is very unstable whenever the ambiguity function of the window is small in magnitude. Consequently, we need to understand when $\mathcal{A}h_{n}(x,y)$, the ambiguity function of the $n$-th Hermite window, is small in absolute value to avoid such regions in our algorithm.
@@
In Figures 2d, 2e, and 2f, the white region reflects the set $S_{n}^{\varepsilon}$ with $\varepsilon=0.1$ and $n=1,5,10$, respectively. The stability set $S_{n}^{\varepsilon}$ is determined by the subset of $\mathbb{R}$ on which the $n$-th Laguerre function $\mathfrak{L}_{n}$ has a large absolute value. Hence, better estimates of $S_{n}^{\varepsilon}$ can aid the design of an efficient numerical strategy for phase retrieval by choosing a minimal set of Hermite windows $\{h_{n_{k}}\}$, such that $\bigcup_{k}S_{n_{k}}^{\varepsilon}$ is as large as possible. As shown in Figures 2g, 2h, and 2i, we plot the union of the stability set $\bigcup_{n\in I}S_{n}^{\varepsilon}$ for different sets of indices $I$. The larger the set $I$, the larger the union of the stability sets on which we can stably recover $\mathcal{A}f(x,y)$.
@@ 3.1.1 Propagation of data noise
(9) $\left\|A-A^{\eta}\right\|_{L^{p,1}(\mathbb{R}^{2})}$ $\leq\frac{C_{p}}{\varepsilon}\left\|\sum_{j=1}^{N}\eta_{j}\right\|_{W^{1,1}(\mathbb{R}^{2})},$
where $C_{p}:=\left\|\mathds{1}_{\Omega}/\omega\right\|_{L^{p,1}(\mathbb{R}^{2})}$, $\mathds{1}_{\Omega}$ denotes the characteristic function on $\Omega$ and $\omega(x,y):=1+2\pi i(x+y)$. Furthermore,
@@
$\left\|A-A^{\eta}\right\|_{L^{p,1}(\mathbb{R}^{2})}=\left\|\mathds{1}_{\Omega}\frac{\sum_{j=1}^{N}T\mathcal{F}(F_{j}-F_{j}^{\eta})}{\sum_{j=1}^{N}\mathcal{A}\phi_{j}}\right\|_{L^{p,1}(\mathbb{R}^{2})}\leq\frac{C_{p}}{\varepsilon}\left\|\omega\,T\sum_{j=1}^{N}\mathcal{F}(F_{j}-F_{j}^{\eta})\right\|_{L^{\infty}(\mathbb{R}^{2})},$
@@
$\left\|A-A^{\eta}\right\|_{L^{p,1}(\mathbb{R}^{2})}$ $\leq\frac{C_{p}}{\varepsilon}\left\|\widetilde{\omega}\,\mathcal{F}\left(\sum_{j=1}^{N}\eta_{j}\right)\right\|_{L^{\infty}(\mathbb{R}^{2})}$
$\leq\frac{C_{p}}{\varepsilon}\left\|\mathcal{F}\left(\sum_{j=1}^{N}(\eta_{j}-\frac{\partial}{\partial x}\eta_{j}+\frac{\partial}{\partial y}\eta_{j})\right)\right\|_{L^{\infty}(\mathbb{R}^{2})}\leq\frac{C_{p}}{\varepsilon}\left\|\sum_{j=1}^{N}\eta_{j}\right\|_{W^{1,1}(\mathbb{R}^{2})},$
with $\widetilde{\omega}:=T\omega=(1+2\pi i(y-x))$ and by employing the Hausdorff-Young inequality in the last step. Finally, (10) can be obtained as
@@
where the last inequality is an application of the Hausdorff-Young inequality. To bound the propagation of data noise in the reconstructions $\widetilde{f}$ and $\widetilde{f\,}^{\eta}$, we first define
@@
(12) $\left\|\widetilde{f}-\widetilde{f\,}^{\eta}\right\|_{L^{p}(\mathbb{R})}$ $\leq\left(\frac{\left\|A\right\|_{L^{p,1}(\mathbb{R}^{2})}}{2(k\,k^{\eta})^{3/2}}+\frac{1}{k^{\eta}}\right)\frac{\max(C_{\mathbb{R},\Omega},C_{p})}{\varepsilon}\left\|\sum_{j=1}^{N}\eta_{j}\right\|_{W^{1,1}(\mathbb{R}^{2})}.$

### v2

@@ 1 Introduction
Existing phase retrieval algorithms are much more involved than the simple inversion formula (3): the most prominent algorithms employed for a wide range of phase retrieval problems are iterative projection algorithms, such as the Gerchberg–Saxton algorithm [15] and Fienup’s algorithm [14], Averaged Alternating Reflections [7] and its variants [11, 13, 21, 20, 10]. A more recent popular iterative scheme is Wirtinger Flow [8]. There also exist methods based on convex relaxation, such as PhaseLift [9], which is based on the idea of lifting the problem to higher dimensions and then recovering a low-rank solution of a linear problem instead of solving the non-linear phase retrieval problem. In practice, however, these algorithms have not been adopted as the iterative methods mentioned before are computationally much more efficient. Another method, specific to STFT phase retrieval, is phase gradient heap integration (PGHI), which recovers the phase of $V_{g}f$ point by point along a grid on the time-frequency plane [23]. By far, we have not described the entire literature on phase retrieval algorithms here. For a thorough description, we refer to [12].
In contrast, the simple formula (3) is not employed in practice because it is generally very unstable: typically, one works with time-frequency localized windows because then, the STFT becomes a time-frequency localized signal representation, i.e., it can better describe which frequencies occurred (roughly) at what time (with the limit being the uncertainty principle). At the same time, fast decay of the window $g$ in time or frequency results in fast decay of $\mathcal{A}g$ in the $x$- or $y$-direction, respectively. Hence, even if $\mathcal{A}g$ is nonzero everywhere on $\mathbb{R}^{2}$, its magnitude typically decays very fast, making recovery of $\mathcal{A}f$ via Equation (3) highly unstable. Take for example the Gauss window $\varphi$. Its ambiguity function is a two-dimensional Gauss function, i.e.,
@@
It is important to emphasize that not only the specific reconstruction formula (3) exhibits poor stability properties when used with a single window function, but Gabor phase retrieval itself (from a single window function) is inherently unstable [4]. The stability properties can be linked to the connectedness of the Gabor transform magnitude as a function in the time-frequency plane [18]. Stability can be restored when global recovery of the phase is relaxed and one only aims at recovering the phase on each connected component individually [3]. In this work, by drawing measurements from several different windows, we overcome the trade-off between stability and disconnectedness of the measurements. We aim at stably recovering the signal globally, instead of on each connected component individually (the latter being inherent to the single window setup). The central question we address in this work is the following:
@@
The rest of the paper is organized to address this key question. In Section 2, we propose two multi-window algorithms for STFT phase retrieval — one utilizing fractional Fourier transformed dilated Gaussian windows, and the other employing Hermite functions of varying degrees. Section 3 establishes mathematical results on the improved inverse problem stability provided by these two multi-window approaches. In Section 4, we present several numerical examples, and conclusions are drawn in Section 5.
@@ 2 Multi-window approaches for STFT phase retrieval
In particular, we propose two architectures for $\left\{g_{j}\right\}_{j=1}^{N}$. The first is based on fractional Fourier transforms of dilated Gaussians, while the second is based on Hermite functions of different order. These choices are based on properties of the ambiguity functions of these windows, but of course, other window families might also be suitable. In the choices we are proposing, the ambiguity functions of the windows are known explicitly and can be related to special functions. This facilitates the analysis as the ambiguity functions of the windows are guaranteed to be bounded away from zero in certain regions. Alternatively, one could numerically estimate the ambiguity functions to derive different sets of window families that are well suited for multi-window STFT phase retrieval.
Before we describe the two measurement systems, we first record some preliminary facts that will be employed.
@@ 2.1 Dilated Gauss windows under the fractional Fourier transform
for some fixed dilation factor $a>1$ and some discretization parameter $N$ such that $\{\alpha_{j}\}_{j=1}^{N}$ is an equidistant discretization of the interval $[0,2\pi]$. In view of Equation (5), this choice of windows results in ambiguity functions $\mathcal{A}\phi_{j}$ that have an elliptical exponential decay stretched by a factor $a$ along the axis $(\cos\alpha_{j},\sin\alpha_{j}).$ The superposition of these “essential supports” is shaped in the form of a daffodil and approaches a disc shape by taking $N$ larger, which is also illustrated in Fig. 1, where all plots are displayed on a logarithmic scale.
@@
Figure 1: The absolute value of (the sum of) ambiguity functions for fractional Fourier transformed (dilated) Gauss windows. All plots are displayed on log-scale. (a) $\mathcal{A}\varphi^{1}$, where $\varphi^{1}$ is the standard Gauss; (b) $\mathcal{A}\varphi^{2}$ for the dilated Gauss with $a=2$; (c) $\sum_{j=1}^{N}\mathcal{A}(\mathcal{F}_{\tilde{\alpha}_{j}}\varphi^{2})$ with $N=8$; (d) $\mathcal{A}\varphi^{10}$ for the dilated Gauss with $a=10$; (e) $\sum_{j=1}^{N}\mathcal{A}(\mathcal{F}_{{\alpha}_{j}}\varphi^{10})$ with $N=4$; (f) $\sum_{j=1}^{N}\mathcal{A}(\mathcal{F}_{{\alpha}_{j}}\varphi^{10})$ with $N=40$.
@@
Remark 2.1.
For given parameters $a,\varepsilon$ and $N$, one can quantify the largest disc that can be covered by the rotated ellipses, therefore guaranteeing that $\mathcal{A}f$ will be recovered on a disc $B_{R}(0)$ for some $R.$ More precisely, this radius is determined by the points of intersection between the ellipses. For this, it suffices to consider a point that lies on the ellipse parametrized by
(6) $ax^{2}+\frac{y^{2}}{a}=\frac{2}{\pi}\left|\ln(\sqrt{2}\varepsilon)\right|$
and its rotated version with rotation angle $\pi/N.$ One point of intersection is then given by intersecting the original ellipse with the line parametrized by $(r\cos(\frac{\pi}{2}-\frac{\pi}{2N}),r\sin(\frac{\pi}{2}-\frac{\pi}{2N}))=(r\sin(\frac{\pi}{2N}),r\cos(\frac{\pi}{2N}))$. Plugging this into (6) results in the radius
$R_{1}^{2}=\frac{C}{D},$
where $C=\frac{2}{\pi}\Bigl{\lvert}\ln\bigl{(}\sqrt{2}\,\varepsilon\bigr{)}\Bigr{\rvert}$, $D=\frac{1}{a}\cos^{2}\!\left(\tfrac{\pi}{2N}\right)+a\,\sin^{2}\!\left(\tfrac{\pi}{2N}\right)$. A corresponding point of intersection is given by
$(\widetilde{x},\widetilde{y})=\left(\sqrt{\frac{C}{D}}\sin{\frac{\pi}{2N}},\sqrt{\frac{C}{D}}\cos\frac{\pi}{2N}\right).$
If, instead, one considers the other pair of intersecting points, one obtains the radius
$R_{2}=\sqrt{\frac{C}{a\,\cos^{2}\!\Bigl{(}\tfrac{\pi}{2N}\Bigr{)}+\frac{\sin^{2}\!\bigl{(}\tfrac{\pi}{2N}\bigr{)}}{a}}}.$
Assuming $a>1$ and $N\geq 3$, we have $R_{1}>R_{2}$. The union of the set of ellipses covers the larger disc. By symmetry, the ambiguity function $\mathcal{A}f$ is then guaranteed to be recovered on $B_{R_{1}}(0)$.
The daffodil-shaped geometry shown in Fig. 1e is a very simple construction and the reconstruction can be further simplified as follows. A simple calculation yields that the ambiguity functions of $\phi_{j}$ are real-valued and positive. Here, we use the notation $R_{\alpha_{j}}(x,y)=(x(\alpha_{j}),y(\alpha_{j}))$:
@@
through pointwise division in (7). We propose Algorithm 1 for the multi-window inversion using fractional Fourier transformed Gaussian windows at various angles.
Algorithm 1 A Multiscale STFT Phase Retrieval Algorithm Based on the Fractional Fourier Transformed Gaussian Windows.
@@
A drawback of this setup is that many windows are required to achieve a good support coverage of the time-frequency plane by their ambiguity functions, such as the disc shown in Fig. 1f. Next, we propose a more technical construction, but with the advantage of requiring less data in the sense that fewer windows are necessary for a good support coverage.
@@ 2.2 Hermite windows
For this second architecture, the set of windows will consist of Hermite (also called Hermite–Gaussian) functions of increasing order. Their ambiguity functions have rotational symmetry with a number of circular root sets related to the order of the Hermite function; see Fig. 2a, Fig. 2b and Fig. 2c. The idea of our approach is to recover $\mathcal{A}f$ as if one was “peeling an onion”: for each Hermite function, we define an annulus on which $\mathcal{A}f$ can be retrieved stably. The union of these annuli results in a large disc in the time-frequency plane. In what follows, we will make this procedure more precise. For this, we start with the definition of the Hermite function $h_{n}$ of order $n$ [1]:
@@
The set of Laguerre functions $\{\mathfrak{L}_{n}(z)\}_{n\in\mathbb{N}}$ forms an orthogonal basis of $L^{2}(\mathbb{R})$ and can be used to express the ambiguity function in (8) as
@@
The $n-$th Laguerre polynomial, and hence also the $n-$th Laguerre function, has exactly $n$ (strictly positive, pairwise different) roots. By Equation (8), $\mathcal{A}h_{n}$ is rotation-invariant. Hence, the root set of $\mathcal{A}h_{n}$ consists of $n$ disjoint circles, centered at the origin $(0,0)$ and thus has Lebesgue measure zero. For stable phase retrieval, we need to determine regions on which $\mathcal{A}h_{n}(x,y)$ is lower bounded in absolute value.
@@
In Fig. 2d, Fig. 2e and Fig. 2f, the white region reflects the set $S_{n}^{\varepsilon}$ with $\varepsilon=0.1$ and $n=1,5,10$, respectively. The stability set $S_{n}^{\varepsilon}$ is determined by the subset of $\mathbb{R}$ on which the $n$-th Laguerre function $\mathfrak{L}_{n}$ has a large absolute value. Hence, better estimates of $S_{n}^{\varepsilon}$ can aid the design of an efficient numerical strategy for phase retrieval by choosing a minimal set of Hermite windows $\{h_{n_{k}}\}$, such that $\bigcup_{k}S_{n_{k}}^{\varepsilon}$ is as large as possible. As shown in Fig. 2g, Fig. 2h and Fig. 2i, we plot the union of the stability set $\bigcup_{n\in I}S_{n}^{\varepsilon}$ for different sets of indices $I$. The larger the set $I$, the larger the union of the stability sets on which we can stably recover $\mathcal{A}f(x,y)$.
@@ 3.1.1 Propagation of data noise
(10) $\left\|A-A^{\eta}\right\|_{L^{p,1}(\mathbb{R}^{2})}$ $\leq\frac{C_{p}}{\varepsilon}\left\|\sum_{j=1}^{N}\eta_{j}\right\|_{{L^{1}}(\mathbb{R}^{2})},$
where $C_{p}:=\left\|\mathds{1}_{\Omega}\right\|_{L^{p,1}(\mathbb{R}^{2})}<\infty$ and $\mathds{1}_{\Omega}$ denotes the characteristic function on $\Omega$. Furthermore,
@@
$\left\|A-A^{\eta}\right\|_{L^{p,1}(\mathbb{R}^{2})}=\left\|\mathds{1}_{\Omega}\frac{\sum_{j=1}^{N}T\mathcal{F}(F_{j}-F_{j}^{\eta})}{\sum_{j=1}^{N}\mathcal{A}\phi_{j}}\right\|_{L^{p,1}(\mathbb{R}^{2})}\leq\frac{C_{p}}{\varepsilon}\left\|\,T\sum_{j=1}^{N}\mathcal{F}(F_{j}-F_{j}^{\eta})\right\|_{L^{\infty}(\mathbb{R}^{2})},$
@@
$\left\|A-A^{\eta}\right\|_{L^{p,1}(\mathbb{R}^{2})}$ $\leq\frac{C_{p}}{\varepsilon}\left\|\mathcal{F}\left(\sum_{j=1}^{N}\eta_{j}\right)\right\|_{L^{\infty}(\mathbb{R}^{2})}\leq\frac{C_{p}}{\varepsilon}\left\|\sum_{j=1}^{N}\eta_{j}\right\|_{{L^{1}}(\mathbb{R}^{2})},$
through employing the Hausdorff–Young inequality in the last step. Finally, (11) can be obtained as
@@
where the last inequality is an application of the Hausdorff–Young inequality.
To bound the propagation of data noise in the reconstructions $\widetilde{f}$ and $\widetilde{f\,}^{\eta}$, we first define
@@
(13) $\left\|\widetilde{f}-\widetilde{f\,}^{\eta}\right\|_{L^{p}(\mathbb{R})}$ $\leq\left(\frac{\left\|A\right\|_{L^{p,1}(\mathbb{R}^{2})}}{2(k\,k^{\eta})^{3/2}}+\frac{1}{k^{\eta}}\right)\frac{\max(C_{\mathbb{R},\Omega},C_{p})}{\varepsilon}\left\|\sum_{j=1}^{N}\eta_{j}\right\|_{{L^{1}}(\mathbb{R}^{2})}.$